# V-HOP: Visuo-Haptic 6D Object Pose Tracking

### v1

@@ Abstract
Humans naturally integrate vision and haptics for robust object perception during manipulation. The loss of either modality significantly degrades performance. Inspired by this multisensory integration, prior object pose estimation research has attempted to combine visual and haptic/tactile feedback. Although these works demonstrate improvements in controlled environments or synthetic datasets, they often underperform vision-only approaches in real-world settings due to poor generalization across diverse grippers, sensor layouts, or sim-to-real environments. Furthermore, they typically estimate the object pose for each frame independently, resulting in less coherent tracking over sequences in real-world deployments. To address these limitations, we introduce a novel unified haptic representation that effectively handles multiple gripper embodiments. Building on this representation, we introduce a new visuo-haptic transformer-based object pose tracker that seamlessly integrates visual and haptic input. We validate our framework in our dataset and the Feelsight dataset, demonstrating significant performance improvement on challenging sequences. Notably, our method achieves superior generalization and robustness across novel embodiments, objects, and sensor types (both taxel-based and vision-based tactile sensors). In real-world experiments, we demonstrate that our approach outperforms state-of-the-art visual trackers by a large margin. We further show that we can achieve precise manipulation tasks by incorporating our real-time object tracking result into motion plans, underscoring the advantages of visuo-haptic perception. Our model and dataset will be made open source upon acceptance of the paper. Project website https://lhy.xyz/projects/v-hop/
1 Introduction
Accurately tracking object poses is a core capability for robotic manipulation, and would enable contact-rich and dexterous manipulations with efficient imitation or reinforcement learning Wen et al. (2022); Li et al. (2024a); Hsu et al. (2024). Recent state-of-the-art object pose estimation methods, such as FoundationPose Wen et al. (2024), have significantly advanced visual tracking by leveraging large-scale datasets. However, relying solely on visual information to perceive objects can be challenging, particularly in contact-rich or in-hand manipulation scenarios involving high occlusion and rapid dynamics.
@@ 1 Introduction
The cognitive science findings show that humans naturally integrate visual and haptic information for robust object perception during manipulation Navarro-Guerrero et al. (2023); Ernst and Banks (2002); Lacey and Sathian (2020). For instance, Gordon et al. (1991) demonstrated that humans use vision to hypothesize object properties and haptics to refine precision grasps. The human “sense of touch” consists of two distinct senses Loomis and Lederman (1986); Dahiya et al. (2010): the cutaneous sense, which detects stimulation on the skin surface, and kinesthesis, which provides information on static and dynamic body posture. This integration, known as haptic perception, allows humans to effectively perceive and manipulate objects Lacey and Sathian (2020). In robotics, analogous capabilities are achieved through tactile sensors (cutaneous sense) and joint sensors (kinesthesis) Navarro-Guerrero et al. (2023).
Drawing inspiration from these human capabilities, researchers have explored the integration of vision and touch in robotics for decades. As early as 1988, Allen (1988) proposed an object recognition system that combined these modalities. More recently, data-driven approaches have emerged to tackle object pose estimation using visuo-tactile information Li et al. (2023); Suresh et al. (2024); Dikhale et al. (2022); Wan et al. (2024); Rezazadeh et al. (2023); Tu et al. (2023); Gao et al. (2023); Li et al. (2024b). Although promising, these methods face two major barriers that hinder their broader applicability: (i) Cross-embodiment: Most approaches overfit specific grippers or tactile sensor layouts, limiting their adaptability. (ii) Domain generalization: Compared to visual-only baselines, visuo-tactile approaches struggle to generalize, hindered by insufficient data diversity and model scalability. Moreover, they typically process each frame independently, which can result in less coherent object pose tracking over sequences in real-world deployments. As a result, existing methods are challenging to deploy broadly and often require significant customization to specific robotic platforms.
@@
We build a multi-embodied dataset with eight grippers using the NVIDIA Isaac Sim simulator for training and evaluation. Compared to FoundationPose Wen et al. (2024), our approach achieves 5% improvement in the accuracy of object pose estimation in terms of ADD-S Xiang et al. (2018) in our dataset. These results highlight the benefit of fusing visual and haptic sensing. In the FeelSight dataset Suresh et al. (2024), we benchmark against NeuralFeels Suresh et al. (2024), an optimization-based visuo-tactile object pose tracker, achieving a 32% improvement in the ADD-S metric and ten times faster run-time speed. Finally, we perform the sim-to-real transfer experiments using Barrett Hands. Our method demonstrates remarkable robustness and significantly outperforms FoundationPose, which could lose object tracks entirely (Fig. 5). When integrated into motion plans, our approach achieves 40% higher average task success rates. To the best of our knowledge, V-HOP is the first data-driven visuo-haptic approach to demonstrate robust generalization across both taxel-based tactile sensors (e.g., Barrett Hand) and vision-based tactile sensors (e.g., DIGIT sensors), as well as on novel embodiments and objects.
@@ 2.1 Problem Definition
In practice, the ground-truth initial pose $\mathbf{T}_{0}$ is hard to obtain and can only be estimated through pose estimation Xiang et al. (2018); Wang et al. (2019a); Park et al. (2019); Li et al. (2024c); Wang et al. (2019b); Lee et al. (2023); Wen et al. (2024); Labbé et al. (2022); He et al. (2022); Liu et al. (2022); Lin et al. (2024a); Tremblay et al. (2018); Wen et al. (2020b). Therefore, we treat $\widehat{\mathbf{T}}_{0}=\mathbf{T}_{0}$ in the following. At each timestep $i$, our model estimates the object pose $\widehat{\mathbf{T}}_{i}$ given all the inputs with the initial pose being the estimate $\widehat{\mathbf{T}}_{i-1}$ at the previous timestep.
The above inputs are the standard inputs from the model-based visual pose tracking problem Wen et al. (2020a); Deng et al. (2021), while the inputs below will serve our haptic representation and will be detailed in later sections.
@@
    Transformation between the camera and the robot frames obtained through hand-eye calibration Marchand et al. (2005).
2.2 Haptic Representation
The effectiveness of haptic learning hinges on its representation. Prior approaches using raw value Lin et al. (2024b), image Guzey et al. (2023), or graph-based Yang et al. (2023); Li et al. (2024b); Rezazadeh et al. (2023) representations often struggle to generalize across diverse embodiments. For instance, Wu et al. (2022) and Guzey et al. (2023) project tactile signals from Xela sensors into a 2D image format. While this allows efficient processing with existing visual models, extending the method to different grippers or sensor layouts proves challenging. Similarly, Li et al. (2024b) and Rezazadeh et al. (2023) employ graph-based mappings, where taxels are represented as vertices. However, variations in sensor layouts result in different graph distributions, creating significant generalization gaps.
In contrast, we adopt a point cloud representation, which naturally encode 3D positions and can flexibly accommodate multi-embodiments. We broadly classify tactile sensors into taxel-based and vision-based. A more comprehensive review on tactile sensors can be found at Yamaguchi and Atkeson (2019). Below, we outline how they are converted into point clouds in prior works Dikhale et al. (2022); Suresh et al. (2024); Watkins-Valls et al. (2019); Falco et al. (2017), paving the way for our unified framework.
@@ 2.2 Haptic Representation
  • Readings ($\mathcal{S}_{r}$): Capturing contact values. Readings are commonly binarized into contact or no-contact states Yin et al. (2023); Xue et al. (2023); Li et al. (2023); Dikhale et al. (2022) based on a threshold $\tau$.
@@
Vision-based sensors. For vision-based tactile sensors Lambeta et al. (2020); Yuan et al. (2017); Donlon et al. (2018); Taylor et al. (2022), the tactile data includes $\mathcal{S}=(\mathcal{S}_{p},\mathcal{S}_{I})$:
@@
  • Images ($\mathcal{S}_{I}$): Capturing contact states using regular RGB image representation. Using the tactile depth estimation model Bauza et al. (2019); Suresh et al. (2024); Kuppuswamy et al. (2020); Suresh et al. (2023, 2022); Ambrus et al. (2021), we can convert $\mathcal{S}_{I}$ into tactile point cloud $\mathcal{S}_{p,c}$.
@@
Yet we are not the first to employ point cloud representations for tactile learning, prior works Dikhale et al. (2022); Suresh et al. (2024); Watkins-Valls et al. (2019); Falco et al. (2017) focus on a single type of sensor and overlook the hand posture. Our key contribution is a unified representation spanning both taxel-based and vision-based sensors on multi-embodiments, empowered by our multi-embodied dataset. We demonstrate generalizability on the Barrett hand (taxel-based) during our real-world experiments and on the Allegro hand (vision-based DIGIT sensor) using the Feelsight dataset Suresh et al. (2024). Our novel haptic representation seamlessly integrates the tactile signals with the hand posture, enabling more effective hand-object interaction reasoning. In subsequent sections, we describe our approach and provide empirical evidence demonstrating that our representation improves generalization capabilities, bridging the gap between heterogeneous tactile sensor modalities.
@@
Figure 2: Network design of V-HOP. The visual modality, based on FoundationPose Wen et al. (2024), uses a visual encoder to process RGB-D observations (real and rendered) into feature maps, which are concatenated and refined through a ResBlock to produce visual embeddings Dosovitskiy et al. (2021). The haptic modality encodes a unified hand-object point cloud, derived from 9D hand $\mathcal{P}_{h}$ and object $\mathcal{P}_{o}$ point clouds, into a haptic embedding that captures hand-object interactions. The red dot in the figure denotes the activated tactile sensor. These visual and haptic embeddings are processed by Transformer encoders to estimate 3D translation and rotation.
3 Methodology
@@
We propose V-HOP, a data-driven approach that fuses visual and haptic modalities to achieve accurate 6D object pose tracking. Our goal is to build a generalizable visuo-haptic pose tracker that accommodates diverse embodiments and objects. We first outline the core representations used in our haptic modality: hand and object representations. Our choice for the representations follows the spirit of the render-and-compare paradigm Li et al. (2018). Later, we introduce our visuo-haptic model and how it is trained. We use the terms “hand” and “gripper” interchangeably.
3.1 Hand Representation
@@
Tactile signals only represent the cutaneous stimulation, while haptic sensing combines tactile and kinesthetic feedback to provide a more comprehensive spatial understanding of contact and manipulation. We propose a novel haptic representation that integrates tactile signals and hand posture in a unified point cloud representation. This hand-centric representation enables efficient reasoning about spatial contact and hand-object interaction.
@@
Using the URDF definition and joint positions $\mathbf{j}$, we generate the hand mesh $\mathcal{M}_{h}$ through forward kinematics and calculate the surface normals. The mesh is then downsampled to produce a 9-D hand point cloud $\mathcal{P}_{h}=\{\mathbf{p}_{i}\}_{i=1}^{n_{h}}$:
@@
  • $[1,0,0]$: Hand point in contact.
  • $[0,1,0]$: Hand point not in contact.
@@
To obtain the contact state of each point, we map the tactile point cloud $\mathcal{S}_{p,c}$, representing the contact points detected by the tactile sensors (Sec. 2.2), onto the downsampled hand point cloud $\mathcal{P}_{h}$. Specifically, for each point in $\mathcal{S}_{p,c}$, we find its neighboring points in $\mathcal{P}_{h}$ within a radius $r$. These neighboring points are labeled as “in contact”, while all others are labeled as “not in contact”. The choice of the radius $r$ is randomized during training and determined by the measured effective radius of each taxel during robot deployment. The resulting haptic point cloud, $\mathcal{P}_{h}$, serves as a unified representation for both tactile and kinesthetic data (Fig. 2).
3.2 Object Representation
We denote the object model point cloud as $\mathcal{P}_{\Phi}=\{\mathbf{q}_{i}\}_{i=1}^{n_{o}}$. Similar to the hand point cloud, $\mathbf{q}_{i}$ follows the same 9-D definitions (Equation 3),
@@ 3.2 Object Representation
To enable reasoning about hand-object interactions, we fuse the hand point cloud $\mathcal{P}_{h}$ and the hypothesized object point cloud $\mathcal{P}_{o}$ to create a hand-object point cloud $\mathcal{P}$,
@@
This novel unified representation adopts the principles of the render-and-compare paradigm from visual approaches Li et al. (2018); Wen et al. (2020a); Labbé et al. (2022); Wen et al. (2024); Tremblay et al. (2023), in which the rendered image (based on pose hypothesis) is compared against the visual observation. The hypothesized object point cloud $\mathcal{P}_{o}$ serves as the “rendered” pose hypothesis (Fig. 2). The hand point cloud $\mathcal{P}_{h}$ represents the real observation using haptic feedback, which we used to compare with. By leveraging this representation, the model captures the contact-rich interactions between the hand and the object by learning feasible object poses informed by haptic feedback.
3.3 Network Design
@@
Visual modality. Unlike prior works, which train the whole visuo-haptic network from scratch, our approach can effectively leverage the pretrained visual foundation model. Our design extends the formulation of FoundationPose Wen et al. (2024), as it demonstrates great generalizability on unseen objects and a narrow sim-to-real gap. To harness the high-quality visual prior captured by it, we utilize its visual encoder $f_{v}$ and freeze it during our training. Using this encoder, We transform the RGB-D observation into visual embeddings $\mathbf{Z}_{v}=f_{v}(\mathbf{O})$.
@@
Haptic modality. In parallel, we encode the hand-object point cloud $\mathcal{P}$ using a haptic encoder $f_{h}$, resulting in a haptic embedding ${\mathbf{Z}_{h}=f_{h}(\mathcal{P})}$. By representing all interactions in point cloud space, our novel haptic representation provides the flexibility to utilize any point cloud-based network for encoding. For this purpose, we choose PointNet++ Qi et al. (2017) as our haptic encoder $f_{h}$. To improve learning efficiency, we canonicalize the point cloud using the centroid of the hand points, ensuring $\mathcal{P}$ is spatially centered around the gripper during processing.
@@
Visuo-haptic fusion. Integrating visual and haptic modalities, however, poses significant challenges. Existing methods often apply fixed or biased weightings between these modalities Li et al. (2023); Suresh et al. (2024); Dikhale et al. (2022); Tu et al. (2023), which can fail under specific conditions. For example, when contact is absent, the visual modality alone should be leveraged, or when occlusion is severe, haptics should be favored. Inspired by the principle of “optimal integration” in human multisensory perception Ernst and Banks (2002); Helbig and Ernst (2007); Lacey and Sathian (2020); Takahashi and Watt (2014); Helbig et al. (2012), where the brain dynamically adjusts the weighting of visual and haptic inputs to maximize reliability, we adopt self-attention mechanisms Vaswani et al. (2017) for the adaptive fusion of visual and haptic embeddings. This ensures robustness across varying scenarios, whether the object is in contact or in clear view.
@@
To achieve this fusion, we propose haptic instruction-tuning, inspired by visual instruction-tuning Liu et al. (2023). While keeping the visual encoder $f_{v}$ frozen, we feed both visual embedding $\mathbf{Z}_{v}$ and haptic embedding $\mathbf{Z}_{h}$ into the original visual-only Transformer encoders Vaswani et al. (2017); Wen et al. (2024), which are initialized with the pretrained weights from FoundationPose. We then fine-tune the Transformer encoders and the haptic encoder $f_{h}$ together. Consequently, visual and haptic information is fused adaptively using self-attention blocks, and the model dynamically adjusts the modality weight based on the context (Fig. 9). Following FoundationPose, we disentangle the 6D pose into 3D translation and 3D rotation and estimate them using two output heads (Fig. 2), respectively.
@@
3.4 Training Paradigm
We train our model by adding noise $(\mathbf{R}_{\epsilon},\mathbf{t}_{\epsilon})$ to the ground-truth pose $\mathbf{T}=(\mathbf{R},\mathbf{t})$ to create the hypothesis pose $\widetilde{\mathbf{T}}=(\widetilde{\mathbf{R}},\widetilde{\mathbf{t}})$:

### v2

@@ Abstract
Humans naturally integrate vision and haptics for robust object perception during manipulation. The loss of either modality significantly degrades performance. Inspired by this multisensory integration, prior object pose estimation research has attempted to combine visual and haptic/tactile feedback. Although these works demonstrate improvements in controlled environments or synthetic datasets, they often underperform vision-only approaches in real-world settings due to poor generalization across diverse grippers, sensor layouts, or sim-to-real environments. Furthermore, they typically estimate the object pose for each frame independently, resulting in less coherent tracking over sequences in real-world deployments. To address these limitations, we introduce a novel unified haptic representation that effectively handles multiple gripper embodiments. Building on this representation, we introduce a new visuo-haptic transformer-based object pose tracker that seamlessly integrates visual and haptic input. We validate our framework in our dataset and the Feelsight dataset, demonstrating significant performance improvement on challenging sequences. Notably, our method achieves superior generalization and robustness across novel embodiments, objects, and sensor types (both taxel-based and vision-based tactile sensors). In real-world experiments, we demonstrate that our approach outperforms state-of-the-art visual trackers by a large margin. We further show that we can achieve precise manipulation tasks by incorporating our real-time object tracking result into motion plans, underscoring the advantages of visuo-haptic perception. Project website: https://ivl.cs.brown.edu/research/v-hop.
I Introduction
Accurately tracking object poses is a core capability for robotic manipulation, and would enable contact-rich and dexterous manipulations with efficient imitation or reinforcement learning [68, 31, 23]. Recent state-of-the-art object pose estimation methods, such as FoundationPose [70], have significantly advanced visual tracking by leveraging large-scale datasets. However, relying solely on visual information to perceive objects can be challenging, particularly in contact-rich or in-hand manipulation scenarios involving high occlusion and rapid dynamics.
@@ I Introduction
The cognitive science findings show that humans naturally integrate visual and haptic information for robust object perception during manipulation [46, 12, 28]. For instance, Gordon et al. [15] demonstrated that humans use vision to hypothesize object properties and haptics to refine precision grasps. The human “sense of touch” consists of two distinct senses [42, 6]: the cutaneous sense, which detects stimulation on the skin surface, and kinesthesis, which provides information on static and dynamic body posture. This integration, known as haptic perception, allows humans to effectively perceive and manipulate objects [28]. In robotics, analogous capabilities are achieved through tactile sensors (cutaneous sense) and joint sensors (kinesthesis) [46].
Drawing inspiration from these human capabilities, researchers have explored the integration of vision and touch in robotics for decades. As early as 1988, Allen [1] proposed an object recognition system that combined these modalities. More recently, data-driven approaches have emerged to tackle object pose estimation using visuo-tactile information [32, 54, 8, 61, 50, 59, 14, 33]. Although promising, these methods face two major barriers that hinder their broader applicability: (i) Cross-embodiment: Most approaches overfit specific grippers or tactile sensor layouts, limiting their adaptability. (ii) Domain generalization: Compared to visual-only baselines, visuo-tactile approaches struggle to generalize, hindered by insufficient data diversity and model scalability. Moreover, they typically process each frame independently, which can result in less coherent object pose tracking over sequences in real-world deployments. As a result, existing methods are challenging to deploy broadly and often require significant customization to specific robotic platforms.
@@
We build a multi-embodied dataset with eight grippers using the NVIDIA Isaac Sim simulator for training and evaluation. Compared to FoundationPose [70], our approach achieves 5% improvement in the accuracy of object pose estimation in terms of ADD-S [72] in our dataset. These results highlight the benefit of fusing visual and haptic sensing. In the FeelSight dataset [54], we benchmark against NeuralFeels [54], an optimization-based visuo-tactile object pose tracker, achieving a 32% improvement in the ADD-S metric and ten times faster run-time speed. Finally, we perform the sim-to-real transfer experiments using Barrett Hands. Our method demonstrates remarkable robustness and significantly outperforms FoundationPose, which could lose object tracks entirely (Fig. 5). When integrated into motion plans, our approach achieves 40% higher average task success rates. To the best of our knowledge, V-HOP is the first data-driven visuo-haptic approach to demonstrate robust generalization across both taxel-based tactile sensors (e.g., Barrett Hand) and vision-based tactile sensors (e.g., DIGIT sensors), as well as on novel embodiments and objects.
@@ II-A Problem Definition
In practice, the ground-truth initial pose $\mathbf{T}_{0}$ is hard to obtain and can only be estimated through pose estimation [72, 62, 48, 36, 63, 30, 70, 27, 18, 40, 37, 57, 67]. Therefore, we treat $\widehat{\mathbf{T}}_{0}=\mathbf{T}_{0}$ in the following. At each timestep $i$, our model estimates the object pose $\widehat{\mathbf{T}}_{i}$ given all the inputs with the initial pose being the estimate $\widehat{\mathbf{T}}_{i-1}$ at the previous timestep.
The above inputs are the standard inputs from the model-based visual pose tracking problem [66, 7], while the inputs below will serve our haptic representation and will be detailed in later sections.
@@
    Transformation between the camera and the robot frames obtained through hand-eye calibration [44].
II-B Haptic Representation
The effectiveness of haptic learning hinges on its representation. Prior approaches using raw value [38], image [16], or graph-based [75, 33, 50] representations often struggle to generalize across diverse embodiments. For instance, Wu et al. [71] and Guzey et al. [16] project tactile signals from Xela sensors into a 2D image format. While this allows efficient processing with existing visual models, extending the method to different grippers or sensor layouts proves challenging. Similarly, Li et al. [33] and Rezazadeh et al. [50] employ graph-based mappings, where taxels are represented as vertices. However, variations in sensor layouts result in different graph distributions, creating significant generalization gaps.
In contrast, we adopt a point cloud representation, which naturally encode 3D positions and can flexibly accommodate multi-embodiments. We broadly classify tactile sensors into taxel-based and vision-based. A more comprehensive review on tactile sensors can be found at [74]. Below, we outline how they are converted into point clouds in prior works [8, 54, 64, 13], paving the way for our unified framework.
@@ II-B Haptic Representation
  • Readings ($\mathcal{S}_{r}$): Capturing contact values. Readings are commonly binarized into contact or no-contact states [78, 73, 32, 8, 34] based on a threshold $\tau$.
@@
Vision-based sensors. For vision-based tactile sensors [29, 79, 10, 56], the tactile data includes $\mathcal{S}=(\mathcal{S}_{p},\mathcal{S}_{I})$:
@@
  • Images ($\mathcal{S}_{I}$): Capturing contact states using regular RGB image representation. Using the tactile depth estimation model [3, 54, 26, 53, 52, 2], we can convert $\mathcal{S}_{I}$ into tactile point cloud $\mathcal{S}_{p,c}$.
@@
Yet we are not the first to employ point cloud representations for tactile learning, prior works [8, 54, 64, 13] focus on a single type of sensor and overlook the gripper posture. Our key contribution is a unified representation spanning both taxel-based and vision-based sensors on multi-embodiments, empowered by our multi-embodied dataset. We demonstrate generalizability on the Barrett hand (taxel-based) during our real-world experiments and on the Allegro hand (vision-based DIGIT sensor) using the Feelsight dataset [54]. Our novel haptic representation seamlessly integrates the tactile signals with the gripper posture, enabling more effective gripper-object interaction reasoning. In subsequent sections, we describe our approach and provide empirical evidence demonstrating that our representation improves generalization capabilities, bridging the gap between heterogeneous tactile sensor modalities.
@@
Figure 2: Network design of V-HOP. The visual modality, based on FoundationPose [70], uses a visual encoder to process RGB-D observations (real and rendered) into feature maps, which are concatenated and refined through a ResBlock to produce visual embeddings [11]. The haptic modality encodes a unified gripper-object point cloud, derived from 9D gripper $\mathcal{P}_{h}$ and object $\mathcal{P}_{o}$ point clouds, into a haptic embedding that captures gripper-object interactions. The red dot in the figure denotes the activated tactile sensor. These visual and haptic embeddings are processed by Transformer encoders to estimate 3D translation and rotation.
III Methodology
@@
We propose V-HOP, a data-driven approach that fuses visual and haptic modalities to achieve accurate 6D object pose tracking. Our goal is to build a generalizable visuo-haptic pose tracker that accommodates diverse embodiments and objects. We first outline the core representations used in our haptic modality: gripper and object representations. Our choice for the representations follows the spirit of the render-and-compare paradigm [35]. Later, we introduce our visuo-haptic model and how it is trained.
III-A Gripper Representation
@@
Tactile signals only represent the cutaneous stimulation, while haptic sensing combines tactile and kinesthetic feedback to provide a more comprehensive spatial understanding of contact and manipulation. We propose a novel haptic representation that integrates tactile signals and gripper posture in a unified point cloud representation. This gripper-centric representation enables efficient reasoning about spatial contact and gripper-object interaction.
@@
Using the URDF definition and joint positions $\mathbf{j}$, we generate the gripper mesh $\mathcal{M}_{h}$ through forward kinematics and calculate the surface normals. The mesh is then downsampled to produce a 9-D gripper point cloud $\mathcal{P}_{h}=\{\mathbf{p}_{i}\}_{i=1}^{n_{h}}$:
@@
  • $[1,0,0]$: Gripper point in contact.
  • $[0,1,0]$: Gripper point not in contact.
@@
To obtain the contact state of each point, we map the tactile point cloud $\mathcal{S}_{p,c}$, representing the contact points detected by the tactile sensors (Sec. II-B), onto the downsampled gripper point cloud $\mathcal{P}_{h}$. Specifically, for each point in $\mathcal{S}_{p,c}$, we find its neighboring points in $\mathcal{P}_{h}$ within a radius $r$. These neighboring points are labeled as “in contact”, while all others are labeled as “not in contact”. The choice of the radius $r$ is randomized during training and determined by the measured effective radius of each taxel during robot deployment. The resulting haptic point cloud, $\mathcal{P}_{h}$, serves as a unified representation for both tactile and kinesthetic data (Fig. 2).
III-B Object Representation
We denote the object model point cloud as $\mathcal{P}_{\Phi}=\{\mathbf{q}_{i}\}_{i=1}^{n_{o}}$. Similar to the gripper point cloud, $\mathbf{q}_{i}$ follows the same 9-D definitions (Equation 3),
@@ III-B Object Representation
To enable reasoning about gripper-object interactions, we fuse the gripper point cloud $\mathcal{P}_{h}$ and the hypothesized object point cloud $\mathcal{P}_{o}$ to create a gripper-object point cloud $\mathcal{P}$,
@@
This novel unified representation adopts the principles of the render-and-compare paradigm from visual approaches [35, 66, 27, 70, 58], in which the rendered image (based on pose hypothesis) is compared against the visual observation. The hypothesized object point cloud $\mathcal{P}_{o}$ serves as the “rendered” pose hypothesis (Fig. 2). The gripper point cloud $\mathcal{P}_{h}$ represents the real observation using haptic feedback, which we used to compare with. By leveraging this representation, the model captures the contact-rich interactions between the gripper and the object by learning feasible object poses informed by haptic feedback.
III-C Network Design
@@
Visual modality. Unlike prior works, which train the whole visuo-haptic network from scratch, our approach can effectively leverage the pretrained visual foundation model. Our design extends the formulation of FoundationPose [70], as it demonstrates great generalizability on unseen objects and a narrow sim-to-real gap. To harness the high-quality visual prior captured by it, we utilize its visual encoder $f_{v}$ and freeze it during our training. Using this encoder, We transform the RGB-D observation into visual embeddings $\mathbf{Z}_{v}=f_{v}(\mathbf{O})$.
@@
Haptic modality. In parallel, we encode the gripper-object point cloud $\mathcal{P}$ using a haptic encoder $f_{h}$, resulting in a haptic embedding $\mathbf{Z}_{h}=f_{h}(\mathcal{P})$}. By representing all interactions in point cloud space, our novel haptic representation provides the flexibility to utilize any point cloud-based network for encoding. For this purpose, we choose PointNet++ [49] as our haptic encoder $f_{h}$. To improve learning efficiency, we canonicalize the point cloud using the centroid of the gripper points, ensuring $\mathcal{P}$ is spatially centered around the gripper during processing.
@@
Visuo-haptic fusion. Integrating visual and haptic modalities, however, poses significant challenges. Existing methods often apply fixed or biased weightings between these modalities [32, 54, 8, 59], which can fail under specific conditions. For example, when contact is absent, the visual modality alone should be leveraged, or when occlusion is severe, haptics should be favored. Inspired by the principle of “optimal integration” in human multisensory perception [12, 19, 28, 55, 20], where the brain dynamically adjusts the weighting of visual and haptic inputs to maximize reliability, we adopt self-attention mechanisms [60] for the adaptive fusion of visual and haptic embeddings. This ensures robustness across varying scenarios, whether the object is in contact or in clear view.
@@
To achieve this fusion, we propose haptic instruction-tuning, inspired by visual instruction-tuning [39]. While keeping the visual encoder $f_{v}$ frozen, we feed both visual embedding $\mathbf{Z}_{v}$ and haptic embedding $\mathbf{Z}_{h}$ into the original visual-only Transformer encoders [60, 70], which are initialized with the pretrained weights from FoundationPose. We then fine-tune the Transformer encoders and the haptic encoder $f_{h}$ together. Consequently, visual and haptic information is fused adaptively using self-attention blocks, and the model dynamically adjusts the modality weight based on the context (Fig. 9). Following FoundationPose, we disentangle the 6D pose into 3D translation and 3D rotation and estimate them using two output heads (Fig. 2), respectively.
@@
III-D Training Paradigm
[…]